OpenStudy (anonymous):

According to the natural sequence of exponents, how would you explain 2-3

OpenStudy (mertsj):

It is -1

OpenStudy (anonymous):

The three is an exponent by the way

OpenStudy (anonymous):

\[2^{-3} ?\]

OpenStudy (anonymous):

Yes:)

OpenStudy (anonymous):

\[2^{-3}\] \[1/2^{3}\] 1/8

OpenStudy (anonymous):

How do you explain that in words though?

OpenStudy (anonymous):

Let we have a base 2 with an exponent power raised to -3 So in order to have a positive exponent factor, we take reciprocal of the number, 1/2^3 Now that the base has a positive exponent number, we simply multiply it the times on the exponent number, so: 1/2*2*2=1/8 is our answer.

OpenStudy (mertsj):

2^4 is 16. Each time the exponent decreases by one, the value of the expression is divided by 2. So 2^3=8, 2^2=4, 2^1=2, 2^0=1, 2^-1=1/2, 2^-2=1/2, and 2^-3=1/8

OpenStudy (mertsj):

\[2^4=16\] \[2^3=8\] \[2^1=2\] \[2^{-1}=\frac{1}{2}\] \[2^{-2}=\frac{1}{4}\] \[2^{-3}=\frac{1}{8}\]

OpenStudy (anonymous):

May be this \[ \huge 2^{-n} = \frac1{2^n} = \prod \limits_{k=1}^ {n} \frac12 \]